# Expanding Chatbot Knowledge in Customer Service: Context-Aware Similar Question Generation Using Large Language Models

### v2

@@ Abstract
Service chatbots play an important role in enhancing customer support by delivering timely responses to diverse queries. Traditionally, these chatbots rely on retrieval-based methods constrained by a predefined knowledge base of question-answer (QA) pairs to guarantee reliable responses. To effectively handle varied customer inquiries, augmenting the knowledge base with similar questions that maintain semantic consistency and linguistic variability is crucial. This paper presents methodologies for a novel approach that utilizes Large Language Models (LLMs) for generating similar questions and selecting an optimal subset of questions for knowledge base augmentation in industrial chatbots. Specifically, we define the SQG task in the context of LLM training and propose a one-to-many objective that incorporates contextual information. We also introduce an optimization framework that selects a diverse subset of similar questions within predefined resource constraints. Experimental results demonstrate significant improvements over traditional methods, achieving greater semantic diversity while aligning with source QA pairs, with over 120% relative improvement in meeting business-specific requirements with human evaluation. Combined with several best practices, we provide a robust, application-driven solution for enhancing chatbot performance and improving customer service satisfaction.
Expanding Chatbot Knowledge in Customer Service: Context-Aware Similar Question Generation Using Large Language Models
Mengze Hong1   Chen Jason Zhang1   Di Jiang2thanks: Corresponding author: dijiang@webank.com   Yuanfeng Song2   Yuanqin He2 Lu Wang3   Zhiyang Su4   Qing Li1 Hong Kong Polytechnic University1  AI Group, WeBank Co., Ltd2 Shenzhen University3  Hong Kong University of Science and Technology4
@@ 1 Introduction
In the era of digital transformation, automated service chatbots have become integral to customer-oriented businesses, particularly in sectors like healthcare and finance, where the accuracy and verifiability of responses are essential Singh et al. (2018); Quamar et al. (2020); Bharadwaj et al. (2017). By leveraging an offline-constructed knowledge base of question-answer (QA) pairs, retrieval-based systems are predominantly employed to ensure reliable responses with zero hallucination. These systems work by matching user queries with pre-existing questions to retrieve answers without involving any generation processes (see Figure 1).
In real scenarios, customer requests can be expressed in various forms, necessitating improved query-matching capabilities Zhang et al. (2025); Wang et al. (2024). Expanding the knowledge base can be carried out offline and seamlessly integrated into production systems, providing great flexibility and computational efficiency. This expansion task is referred to as Similar Question Generation (SQG), where the generated questions need to be possibly diverse so that a broader spectrum of expressions can be captured while maintaining semantic consistency with the source question to uphold the integrity of the question-answer relationship. Traditionally, SQG tasks were conducted via human crowdsourcing due to the necessity of domain-specific knowledge, which was expensive and unable to generate diverse similar questions due to the independency Liu et al. (2022). To overcome this limitation, rule-based methods were adopted to automate this process, notably used in text data augmentation Feng et al. (2021); Sennrich et al. (2016). Moreover, specialized models, such as SimBERT Su (2020) and RoFORMER-Sim Su (2021), have been employed to ensure better semantic consistency of generated text, but their performance in terms of diversity is hindered by the lack of domain-specific knowledge, often resulting in repetitive or over-generic questions.
Figure 1: Simplified schematic overview of a retrieval-based chatbot with a knowledge base that consists of a set of QA pairs. The yellow region highlights the questions augmented by the similar question generation. The incoming customer’s request is matched with the questions in QA pairs, and the answer to the best-matched question is selected as the response to the customer.
Recently, Large Language Models (LLMs) have demonstrated unprecedented natural language understanding (NLU) and generation (NLG) capability Wei et al. (2022); OpenAI (2023); Hong et al. (2025). They can be used for a wide range of downstream tasks via prompting Radford et al. (2019) or fine-tuning Bao et al. (2023). However, repeated generation employing the naive sequence-to-sequence paradigm is ineffective for generating a large number of diverse questions due to the lack of control over the generation process, often resulting in questions that are either similarly expressed or significantly deviate from the meaning of the source question, unfulfilling business requirements.
In this work, we propose a novel approach for expanding the knowledge base of service chatbots with LLMs, aiming at generating a large number of diverse questions while maintaining semantic consistency with the source question. To sum up, our contributions are as follows:
  • We present a pioneering effort in defining the Similar Question Generation task in the context of LLM training and inference. Motivated by the limitation of the conventional “one-to-one” generation, we propose novel methods that employ a context-aware one-to-many training and inference target. This approach is further enhanced by incorporating the QA answer as contextual guidance, which diversifies the generated questions to a broader semantic space and ensures consistency with the underlying intent of the original QA pair.
  • With a production dataset from an active customer service chatbot, we demonstrate the superior performance of the proposed method in generating a substantial number of similar questions with high diversity and semantic consistency. Qualitative evaluations by human experts further confirm the practical advantages of the proposed methods, resulting in over 120% relative improvements in the number of satisfactory questions being generated.
  • We present an optional constrained optimization framework aimed at selecting an optimal subset of similar questions that maximizes diversity while adhering to budget constraints for storage and retrieval, making the proposed methods adaptable to various real-world applications and operational environments.
{CJK}
UTF8gbsn
@@
Table 1: Illustration of conventional paradigm and proposed methods for fine-tuning and inference of LLMs.
@@ 2 Proposed Methods
In this section, we first formulate the setting of similar question generation and then elaborate on our proposed LLM-based methods.
2.1 Similar Question Generation
Similar Question Generation (SQG) aims to create a diverse yet semantically consistent set of questions that can be matched to a specific answer in a knowledge base Gollapalli and Ng (2022). In this context, semantic consistency refers to the preservation of the original intent and meaning, ensuring the generated questions can still be accurately matched to the intended answer in the knowledge base Hong et al. (2024). Conversely, syntactic diversity pertains to the variation in phrasing and structure of the generated questions, allowing for different expressions that are crucial for improved query-matching Guo et al. (2024); Ma et al. (2023). The task of generating similar questions is formally defined as follows:
Definition.
Let $D=\{(q^{i},a^{i})\}_{i=1}^{N}$ represent a predefined knowledge base containing $N$ QA pairs, where each QA pair consists of a question and its associated answer. The goal of SQG is to generate a set of $K$ similar questions $Q^{i*}=\{q^{i}_{1},q^{i}_{2},\ldots,q^{i}_{K}\}$ for each given question $q^{i}$ such that these questions correspond to the same answer $a^{i}$. Thus, the knowledge base is augmented into $D^{*}=\{(\{q_{j}^{i}\}_{j=1}^{K_{i}},a^{i})\}_{i=1}^{N}$.
$\mathcal{L}_{ft}=-$ $\sum_{j}\sum_{i}log(P_{\Phi}(q_{j}|q_{i})).$ (1)
@@ 2.1 Similar Question Generation
To effectively generate diverse similar questions, we propose leveraging the Natural Language Understanding (NLU) capabilities and autoregressive generation of LLMs. By prompting the model to generate multiple similar questions in a single step, we draw inspiration from the concept of batch prompting Cheng et al. (2023) but extend beyond inference efficiency to enhance contextual awareness of the previously generated questions.
2.2 Context-Aware Batch Generation
To enhance control over the generation process, we introduce a novel objective known as the one-to-many paradigm. This method enables the LLM to generate multiple similar questions in response to a single source question. The corresponding “Context-Aware Batch Generation,” approach is presented in the second block of Table 1, which provides greater control over the generation process. During training, the LLM learns to identify semantic similarities and subtle expressive differences among multiple target questions. In the inference phase, the auto-regressive nature of LLMs Vaswani et al. (2017) allows for the incorporation of previously generated questions, which helps regularize subsequent outputs and reduces the likelihood of generating repetitive or excessively divergent questions. This concept is illustrated in the blue region of Figure 2 (b), where the generated questions are spaced apart, facilitating broader coverage of the desired semantic landscape.
@@ 2.3 Intention-Enhanced Batch Generation
While the one-to-many objective enables a more effective exploration of the semantic space surrounding the source question, this space is constrained by the strict requirement for semantic consistency. Since the query-matching process in the retrieval primarily acts as a proxy for selecting the answer that best aligns with the customer’s intent, this motivates the incorporation of the source answer as semantic guidance, enabling greater alignment. Since the SQG is an augmentation process with known question-answer pairs, integrating the answer can also be viewed as introducing contextual prior knowledge into the generation process. The corresponding prompt is presented in the third block of Table 1. When visualized in the semantic space, as shown in the green region of Figure 2 (b), this approach expands the exploration beyond the immediate vicinity of the source question and skews towards the desired answer.
2.4 Refined Training and Inference
Formally, given a set of similar questions, ($q_{1}$, $\dots$, $q_{N}$), and the corresponding answer $a$, we construct training sample as (($q_{i}$, $a$), ($q_{j+1}$, $\dots$, $q_{j+L}$)), where $q_{i}$ and $a$ serve as the input, and the randomly sampled $L$ questions are the prediction target. The loss for predicting each question $q_{j+l}$ is defined as:
@@ 2.4 Refined Training and Inference
where the generation of the question $q_{j+l}$ is conditioned on both the QA pair, $(q_{i},a)$ and the previously generated similar questions, $(q_{j+1},\dots,q_{j+l-1})$. Consequently, the overall training objective can be formulated as:
$\mathcal{L}_{Intention}=$ $\sum_{i}\sum_{j}\sum_{l=1}^{L}\mathcal{L}_{j+l}(q_{i},a).$
3 Experimental Setup
3.1 Dataset
The dataset, sourced from an active customer service chatbot in the banking industry, comprises over 3,000 question-answer pairs in Chinese, with an average of 40 human-annotated similar questions corresponding to each source QA [^1] The full dataset will be released upon acceptance.. From this, we constructed a training dataset of 90,000 instances by randomly sampling the raw QA pairs, following the format outlined in Table 1. For the quantitative evaluation, we utilized 90 unseen QA pairs, each with an average of 45 reference questions. In the human assessment, we collected 15 new questions from the recent records of the service chatbot, reflecting practical use cases.
@@ 3.1 Dataset
Improvement (%) 0.07% 3.06% 1.82% - - - 121.64%
Table 2: Performance comparison of proposed batch generation methods and baseline methods. The semantic relevance and character-level diversity are obtained when generating 20 similar questions. The acceptance ratio is obtained by human evaluation on 15 newly collected questions.
3.2 Evaluation Metrics
Semantic Relevancy.
Given a set of generated questions $q=[q_{1},q_{2},\dots,q_{n}]$ and a set of pre-annotated reference questions $R=[r_{1},r_{2},\dots,r_{m}]$, the precision score for each generated question $q_{i}$ is computed as the maximum BERTScore value Zhang et al. (2020) between $q_{i}$ and the reference set $R$, serving as an indicator of the degree of semantic consistency with reference questions. Conversely, the recall metric is determined through an inverse process and quantifies the extent of semantic diversity among the generated questions. The F1 score is a harmonic mean of precision and recall, providing a single metric that balances the trade-off. A higher F1 score indicates a better balance between precision and recall, suggesting that the generated questions are both relevant and sufficiently diverse to cover a broader range of semantic meanings.
Character-Level Diversity.
In addition, we evaluate the character-level diversity of generated questions with the Distinct-N metric Li et al. (2016), which assesses how unique the generated outputs are by measuring the number of distinct N-grams within the joint outputs. We utilize Distinct-1 and Distinct-2 to measure the lexical diversity of output words, as well as word combinations and transitions, respectively. Furthermore, we calculate their average, Distinct-Avg, to represent the overall character-level diversity. A larger value indicates greater textual diversity in the expression.
Qualitative Evaluation.
The quantitative metrics, while informative, do not account for text fluency and semantic coherence to the corresponding answer Chiang and Lee (2023). To accommodate these aspects, we introduce human evaluations that closely align with real-world business requirements. Specifically, the generated similar questions are examined by five business experts with reference to the source QA, and those that meet the criteria are marked as “accepted”, and the acceptance ratio is calculated as the proportion of accepted questions to the total number of generated questions.
The quantitative metrics calculations and human evaluation criteria are detailed in Appendix A.
4 Experimental results
In this section, we first present the experimental results of the proposed method. Then, we offer additional insights and best practices with ablation studies. Finally, we discuss the use of generated questions with an optimization framework for augmenting the chatbot knowledge base. The implementation details are presented in Appendix B.

### v3

# Augmenting Compliance-Guaranteed Customer Service Chatbots: Context-Aware Knowledge Expansion with Large Language Models

@@ Abstract
Retrieval-based chatbots leverage human-verified Q&A knowledge to deliver accurate, verifiable responses, making them ideal for customer-centric applications where compliance with regulatory and operational standards is critical. To effectively handle diverse customer inquiries, augmenting the knowledge base with “similar questions” that retain semantic meaning while incorporating varied expressions is a cost-effective strategy. In this paper, we introduce the Similar Question Generation (SQG) task for LLM training and inference, proposing context-aware approaches to enable comprehensive semantic exploration and enhanced alignment with source question-answer relationships. We formulate optimization techniques for constructing in-context prompts and selecting an optimal subset of similar questions to expand chatbot knowledge under budget constraints. Both quantitative and human evaluations validate the effectiveness of these methods, achieving a 92% user satisfaction rate in a deployed chatbot system, reflecting an 18% improvement over the unaugmented baseline. These findings highlight the practical benefits of SQG and emphasize the potential of LLMs, not as direct chatbot interfaces, but in supporting non-generative systems for hallucination-free, compliance-guaranteed applications.
Augmenting Compliance-Guaranteed Customer Service Chatbots: Context-Aware Knowledge Expansion with Large Language Models
Mengze Hong1,2, Chen Jason Zhang1, Di Jiang1thanks: Corresponding Author, Yuanqin He2 1Hong Kong Polytechnic University, 2AI Group, WeBank Co., Ltd
Figure 1: Schematic overview of a compliance-guaranteed chatbot with a predefined knowledge base for Match-and-Respond. The yellow region highlights the questions augmented by the similar question generation.
@@ 1 Introduction
Customer service automation is essential for digital transformations, commonly deploying AI-driven chatbots to handle diverse inbound customer inquiries to reduce the workload in labor-intensive call centers and enable timely responses across online platforms Jiang et al. (2025). However, popular generative language models are prone to hallucinations, generating inconsistent or incorrect answers Huang et al. (2025), making LLM-driven chatbot interfaces infeasible in sectors like healthcare and finance, where reliability and verifiability of responses are critical concerns Singh et al. (2018); Bharadwaj et al. (2017). To ensure compliance with regulatory and operational requirements, retrieval-based chatbot systems, a framework established before LLMs, use human-verified question-answer (QA) pairs from an offline knowledge base to deliver hallucination-free responses Wu et al. (2018). As illustrated in Figure 1, these systems employ a Match-and-Respond process, matching input queries to existing questions to retrieve accurate responses, eliminating generation needs.
Customer
我开了证明怎么还没收到?
I applied for the certificate, why haven’t I received it yet?
Source
证明开具时间要多久?
How long does it take to process the certificate?
Generated
开了证明什么时候能收到?
Applied for the certificate, when can I receive?
Table 1: Customer query matching with source question (failed) and generated similar question (success).
In practice, customer queries exhibit high diversity in expression, and a failure in query matching may interrupt the interaction and cause user dissatisfaction Zhang et al. (2025); Wang et al. (2024). Expanding chatbot knowledge offline enables seamless integration into production systems, offering a static augmentation approach to enhance query-matching performance with minimal computational load and runtime latency, ensuring practicality for lightweight industrial deployment. This expansion can be effectively achieved through the Similar Question Generation (SQG) task, where each source question in the predefined knowledge base is augmented with multiple “similar questions” that are diverse in expression while preserving semantic consistency to maintain the question-answer relationship. As demonstrated in Table 1, a similar question more effectively matches the user query.
Traditionally, data augmentation relies on costly human crowdsourcing, offering limited diversity due to its independent nature Liu et al. (2022). Rule-based automation methods, such as SimBERT Su (2020) and RoFORMER-Sim Su (2021), improve semantic consistency but lack contextual awareness, producing repetitive or generic outcomes Feng et al. (2021). The emergence of LLMs highlights their potential for data augmentation through advanced language understanding and generation capabilities Wei et al. (2022); Hong et al. (2025a), improved by prompting Liu et al. (2023) and fine-tuning Bao et al. (2023) techniques, which frequently outperform human workers in cost efficiency and performance Gilardi et al. (2023); Törnberg (2023). However, the SQG task presents a unique challenge in requiring diversity among generated questions, where standard sequence-to-sequence methods struggled to produce varied questions due to limited control over the generation process Jiang and de Rijke (2018), necessitating tailored model training and inference strategies.
In this work, we propose novel approaches for augmenting customer service chatbots with similar questions generated by LLMs, highlighting the importance of contextual awareness in the guided generation process. The contributions include:
  1. 1.
    To the best of our knowledge, we are the first to define the SQG task for retrieval-based service chatbot augmentation, formulating LLM training and inference, and proposing context-aware one-to-many generation paradigms.
  2. 2.
    We present budget-constrained optimization techniques to select prompt demonstrations and similar question subsets, facilitating knowledge base expansion and ensuring cost-effective, adaptable deployment across diverse real-world application scenarios.
  3. 3.
    Experiments demonstrate over 120% relative improvement in qualitative assessment, 4.74% increase in total diversity, and 18% enhancement in user satisfaction compared to unaugmented chatbot systems.
2 Problem Formulation and Background
@@
Table 2: Illustration of conventional generation and proposed methods for fine-tuning and inference of LLMs.
2.1 Problem Formulation
Similar Question Generation aims to create a diverse yet semantically consistent set of questions that can be matched to a specific answer in a knowledge base. In this context, semantic consistency refers to the preservation of the original intent and meaning (e.g., “inquire-promotion”), ensuring the generated questions can still be accurately matched to the correct answer in the knowledge base Gollapalli and Ng (2022); Hong et al. (2025c); Jiang et al. (2021). Conversely, syntactic diversity pertains to the variation in phrasing and structure of the generated questions, enabling different expressions in the knowledge base that are essential for enhancing query-matching Guo et al. (2024); Ma et al. (2023).
2.2 SQG Training and Inference with LLMs
Conventional methods that utilize LLMs for similar question generation typically adhere to a naive sequence-to-sequence approach, referred to as the one-to-one paradigm. In this approach, the LLM generates a single question at a time in response to a given source question. For a set of similar questions $(q_{1},\dots,q_{K})$, we can construct training samples by pairing questions, such as $\{(q_{1},q_{2}),\dots,(q_{1},q_{K})\}$. A typical prompt template is illustrated in the first block of Table 2. Given a generative language model $P_{\Phi}(y|x)$ with parameters $\Phi$, the training objective can be formulated as maximizing the following language modeling objective:
$\mathcal{L}_{ft}=$- $\sum_{j}\sum_{i}log(P_{\Phi}(q_{j}|q_{i})).$ (1)
@@ 3 Proposed Methods
3.1 Context-Aware Batch Generation
To enhance control over the generation process, we introduce the one-to-many paradigm. This method enables the LLM to generate multiple similar questions in response to a single source question (see the second block of Table 2). During training, the LLM learns to identify semantic similarities and subtle expressive differences among multiple target questions. In the inference phase, the auto-regressive nature of LLMs Vaswani et al. (2017) allows for the incorporation of previously generated questions, which helps regularize subsequent outputs and reduces the likelihood of generating repetitive or excessively divergent questions. While one-to-many generation, or Batch Prompting Cheng et al. (2023), is typically used for cost-saving and often delivers lower performance compared to standard prompting, we argue that incorporating previously generated questions into autoregressive generation is highly effective for the SQG task, which introduces contextual guidance and leads to more diversified questions with lower generation cost.
@@ 3.2 Intention-Enhanced Batch Generation
While the one-to-many paradigm enables a more effective exploration of the semantic space surrounding the source question, this space is constrained by the strict requirement for semantic consistency. Since the SQG is an augmentation process with known question-answer pairs, integrating the source answer can also be viewed as introducing contextual prior knowledge into the generation process. The corresponding formulation is presented in the third block of Table 2. When visualized in the semantic space, as shown in the green region of Figure 2 (b), this approach expands the exploration beyond the immediate vicinity of the source question and skews towards the desired answer.
3.3 Refined Training Objective
With the proposed one-to-many generation, we refine the training objective to generate multiple similar questions from a single QA pair. Formally, given a set of similar questions, ($q_{1}$, $\dots$, $q_{N}$), and the corresponding answer $a$, we construct training sample as $($$q_{i}$, $a$), ($q_{j+1}$, $\dots$, $q_{j+L}$)), where $(q_{i},a)$ is the input QA pair, and $(q_{j+1}$, $\dots$, $q_{j+L})$ are $L$ target questions; here, $i$ indexes the input question from the knowledge base ($i=1,\dots,M$), and $j$ iterates over starting indices of $L$ consecutive similar questions ($j=1,\dots,N-L$). The loss for predicting each target question $q_{j+l}$ is defined as:
$\begin{aligned} \mathcal{L}_{j+l}(q_{i},a)=-\log(P_{\Phi}(q_{j+l}|q_{i},a,q_{j+1},\dots,q_{j+l-1})),\end{aligned}$
where the generation of the question $q_{j+l}$ is conditioned on both the original QA pair $(q_{i},a)$ and all previously generated similar questions in the same batch, $(q_{j+1},\dots,q_{j+l-1})$, enabling the model to capture contextual dependencies and improve syntactic diversity while maintaining semantic consistency. Consequently, the overall training objective can be formulated as:
$\mathcal{L}_{Intention}=$ $\sum_{i}\sum_{j}\sum_{l=1}^{L}\mathcal{L}_{j+l}(q_{i},a).$ (2)
4 Optimization Framework
Algorithm 1 Question Subset Mining
1:Knowledge base $\mathcal{D}=\{(q_{i},a_{i})\}_{i=1}^{M}\setminus\{(q_{s},a_{s})\}$, source question $q_{s}$, target $K$
2:Question subset $\mathcal{P}\subseteq\mathcal{D}$
@@ 4 Optimization Framework
4:for $(q_{i},a_{i})\in\mathcal{D}$ do
5:  $\phi(q_{i})\leftarrow\text{S}(q_{s},q_{i})$
6:  $L\leftarrow L\cup\{(q_{i},a_{i},\phi(q_{i}))\}$
7:end for
8:Order $L$ by $\phi(q_{i})$ in descending sequence
[…]
10:while $|\mathcal{P}|<K$ and $L\neq\emptyset$ do
11:  Select $z^{\prime}\in L$ maximizing $\sum_{z\in\mathcal{P}}\text{dist}(q_{z},q_{z^{\prime}})+\phi(q_{z^{\prime}})$
12:  $\mathcal{P}\leftarrow\mathcal{P}\cup\{z^{\prime}\}$, $L\leftarrow L\setminus\{z^{\prime}\}$
13:end while
14:return $\mathcal{P}$
4.1 Dynamic Demonstration Selection for Contextual Prompting
To enhance contextual generation, we construct in-context prompts using the knowledge base $\mathcal{D}={(q_{i},a_{i})}_{i=1}^{M}$, where some questions $q_{i}$ are associated with similar questions $\mathbf{q}_{i}^{*}$. The goal is to select $K$ examples for a target question $q_{s}$ and append them with their similar questions to enable in-context learning. The objective is:
$\begin{aligned} \arg\max_{\mathcal{P}\subseteq\mathcal{D},|\mathcal{P}|=K}\ &\left[\sum_{i=1}^{K}\text{S}(q_{s},q_{p_{i}})+\frac{\alpha}{K}\sum_{i\neq j}\text{dist}(q_{p_{i}},q_{p_{j}})\right],\end{aligned}$
where $\text{S}(q_{s},q_{p_{i}})$ is the cosine similarity between BERT embeddings of $q_{s}$ and $q_{p_{i}}$, ensuring relevance, $\text{dist}(q_{p_{i}},q_{p_{j}})$ is the Euclidean distance between question embeddings, measuring diversity Qian et al. (2004), and $\alpha$ is a tunable constant that normalizes the diversity contribution to approximately linear scaling ($\alpha=0.5$). To solve this optimization problem, we propose the Question Subset Mining (QSM) algorithm, designed to balance relevance and diversity (see Algorithm 1).
Algorithm 2 Greedy Algorithm for Maximizing Pairwise Diversity
2:Subset $S\subseteq Q^{*}$
[…]
4:while $B_{r}>0$ and $Q^{*}\setminus S\neq\emptyset$ do
5:  Select $q^{*}\in Q^{*}\setminus S$ that maximizes:
6:  if $\text{cost}(q^{*})\leq B_{r}$ then
7:   $S\leftarrow S\cup\{q^{*}\}$
8:   $B_{r}\leftarrow B_{r}-\text{cost}(q^{*})$
9:  else
10:   $Q^{*}\leftarrow Q^{*}\setminus\{q^{*}\}$
11:  end if
12:end while
13:return $S$
4.2 Similar Question Selection for Knowledge Base Expansion
Instead of generating a small set of questions directly, we argue that a two-step approach, which generates many candidate questions and then selects the best subset, would offer greater flexibility and achieve better results in industrial applications. The constrained optimization problem is formally defined as:
$\max_{S\subseteq Q^{*}}\sum_{\begin{subarray}{c}q_{a},q_{b}\in S\\ q_{a}\neq q_{b}\end{subarray}}\text{dist}(q_{a},q_{b})\quad\text{s.t.}\quad\sum_{q\in S}\text{cost}(q)\leq B.$ (4)
where $\text{cost}(q)$ denotes the cost of including question $q$, either by token length or a uniform cost.
With the proof of NP-hardness and submodularity presented in Appendix D, we propose a greedy algorithm that efficiently approximates the optimal solution with a guaranteed approximation bound of 11/e1-1/e. The greedy algorithm iteratively selects the sample $q^{*}$ that provides the highest marginal gain relative to its cost while satisfying the budget constraint (see Algorithm 2).
@@ 4.2 Similar Question Selection for Knowledge Base Expansion
+ dynamic demo selection 0.8612 0.8527 0.8569 0.2105 0.3627 0.2866 $82.0\%$
Improvement (%) 0.07% 4.74% 2.60% - - - $121.64\%$
Table 3: Performance comparison of similar question generation methods. The universal best results are bolded, and the best results among baseline methods are underlined to compute relative improvement.
5 Experimental Setup
5.1 Dataset
To evaluate the proposed methods for generating similar questions, we leverage a dataset sourced from an active customer service chatbot in the financial sector, which comprises over 3,000 QA pairs in Chinese, each with an average of 40 human-annotated similar questions. From this, we constructed a training dataset of 90,000 instances by randomly sampling the raw QA pairs, following the format outlined in Table 2. Additional experiments with public datasets are presented in Appendix C for completeness and reproducibility.
5.2 Evaluation Details
For the quantitative evaluation, we utilized 90 unseen QA pairs, each with an average of 45 reference questions. In the human assessment, we collected 15 new questions from the recent records of the service chatbot, reflecting practical use cases. We report the following performance metrics:
Semantic Relevancy
Precision is the maximum BERTScore Zhang et al. (2020) between each generated question and reference question, measuring semantic consistency. Recall is computed inversely, assessing semantic diversity. The F1 score measures the harmonic mean of precision and recall, balancing relevance and diversity.
Character-Level Diversity
We use Distinct-N Li et al. (2016) to evaluate lexical diversity and report the Distinct-1, Distinct-2, and their average, Distinct-Avg score, counting unique N-grams in generated questions. Higher values indicate greater textual diversity.
Qualitative Evaluation
Five industry experts assess generated questions against source QA, marking acceptable ones based on the semantic consistency and syntactic diversity criteria [^1] Metrics and evaluation criteria are detailed in Appendix A. Implementation details are presented in Appendix B..
6 Results and Discussions